Given a second order linear homogeneous differential equation a2(x)y + a1(x)y + 20 (x)y = 0 we know that a fundamental setPlease show all work and steps! Would like to learn how!
Answer #1

Given fowblem is xy +42y-189=0. yt deny = 0 Or, y+ y = 23. pix) = (1/2 and e Skesde usted -4 logee se e e Sants - Yah = 1254-4y/+4y=0 Its characteristic equation is. 25m2-4m+ 4 = 0 me 4 /16-400 - 4/-384 T 50 50 4+856 2 25 = 50 0,- en moderne je